# Inferring resource competition in microbial communities from time series

### v1

@@ II.1 Environment-mediated consumer-resource model
Figure 1: Environment-mediated ecological interactions between resources and consumers. (a) Cartoon of interactions among environment-mediated resources and consumer species. (b) An example of the consumer-resource interaction matrix $\mathbf{C}$ with elements $C_{i\alpha}=r_{i\alpha}\gamma_{i\alpha}$ (colorbar). The probability of non-zero consumer-resource interactions depends on whether the resource is preferred by the species or not, with an adjustable parameter $q$ that encodes guild-structure bias ($q=0.4$ for this example). All resources also have private resources to prevent extinction. (c) Effective interactions among the species are captured by the effective resource-utilization overlap matrix, $\mathbf{C}\mathbf{C}^{\intercal}$ (colorbar), shown here for 3 different values of the guild-structure bias $q$.
@@
The consumer-resource interaction matrix $\mathbf{C}$, where each entry is $C_{i\alpha}=r_{i\alpha}\gamma_{i\alpha}$, describes the intake of resource $\alpha$ by species $i$ (Fig. 1(b)). Summing over the resources we obtain the resource-utilization overlap matrix $\mathbf{C}\mathbf{C}^{\intercal}$ (Fig. 1(c)). Large off-diagonal entries of this matrix ($\mathbf{C}\mathbf{C}^{\intercal}_{i,j}$) indicate that species $i$ and $j$ have high overlap in the resources they utilize. Species with high resource-overlap are expected to compete, and those without are expected not to compete. One can consider the resource-utilization matrix as a proxy for effective interactions between species, which is informative about the community structure.
Specifically, we are interested in detecting guilds among the interacting species. To introduce a guild structure in resource overlap, we write the uptake rate as $r_{i\alpha}=rg_{i\alpha}$, where $g_{i\alpha}$ is an adjacency matrix that takes the value of 0 if species $i$ does not uptake resource $\alpha$, and 1 if it does consume this resource. In order to generate ensembles of random interaction strengths such that the resulting $\mathbf{C}\mathbf{C}^{\intercal}$ matrix has a block (guild) structure, we construct a bipartite graph between $N$ species and $M$ resources. We then separate both the species and the resources into $k$ groups, with each group of species preferring to consume resources from one group of resources. The elements of the adjacency matrix that specifies these preferences, $g_{i\alpha}$, are drawn from a Bernoulli distribution parametrized by a probability $p_{\text{in}}$. For the other $k-1$ groups of non-preferred resources, the elements of the adjacency matrix $g_{i\alpha}$ are drawn from a Bernoulli distribution parametrized by a probability $p_{\text{out}}<p_{\rm in}$. In addition, each species is assigned its private resource to prevent extinction (Fig. 1(b)). The yield $\gamma_{i\alpha}$ is drawn from a Gaussian distribution with a positive mean and small variance to introduce randomness. All yields are non-negative (see Methods for specific parameter values).
@@
The consumer-resource model (1)-(2), environmental drive (5)-(7), and resource guild structure $r_{i\alpha}$ among consumers, completely define the model. To illustrate its behavior and gain intuition, we first simulated a small network with $N=6$ species and $M=30$ resources, with a guild-structure bias of $q=0.4$. The resource-utilization matrix $\mathbf{C}\mathbf{C}^{\intercal}$ is given by Fig. 2(a). The specific set of parameters is given in Methods IV.1. The first $N_{a}=3$ species form one guild, while the remaining $N_{b}=3$ species form the other. To study the impact of environmental drive on species abundances, we simulate the ECRM under sinusoidal and OU environmental drives each sampled at three different ranges of timescales – slow, fast, and with a mixture of timescales (Fig. 2(b)). Example traces of pairs of species within the same guild show complex coupled dynamics depending on the environmental timescales (Fig. 2(d)).
@@ II.2 Pairwise measures of species couplings
To explore how the coupled dynamics of species abundances can be informative about their resource overlap $\mathbf{C}\mathbf{C}^{\intercal}$, we consider observables of increasing complexity that measure the deviation from independence between the time traces of two species.
@@ II.2.1 Equal-time correlation $\mathbf{C}(0)$
Specifically, as shown in Figure 2(c) in our example of 6 species distributed into 2 guilds, when the environmental drive is fast, $C_{ij}(0)\propto\left(\mathbf{C}\mathbf{C}^{\intercal}\right)_{ij}$, while when the environmental drive is slow, $C_{ij}(0)\propto-\left(\mathbf{C}\mathbf{C}^{\intercal}\right)_{ij}$. Intuitively this happens because fast resource dynamics drive coherent responses of members within a guild, but slow resource dynamics lead to competition within guilds (the underlying mechanism is first explored and explained in detail in crocker_microbial_2025 ). This problem is more pronounced when environmental fluctuations are driven by a mixture of timescales (middle panel in Fig. 2(c)). There is no correlation between the equal-time correlation $\mathbf{C}(0)$ and the resource utlization overlap $\mathbf{CC}^{\intercal}$ ($p$-value = 0.4, Spearman correlation test). In this example, the pair of species with the strongest resource utlization overlap (indicated by red boxes in Fig. 2(a)) has a equal-time correlation close to 0, as indicated by the black arrow in Fig. 2(c).
Without knowing a priori the timescale of the environmental fluctuation, the equal-time correlation $\mathbf{C}(0)$ does not return a good enough estimator of the resource utlization overlap.
@@ II.2.2 Cross-correlation with time-delay reveals species interaction structure
Since equal-time correlations do not reveal the stucture of effective species-species interactions, we took a closer look at time series data. Let us focus again on the pair of species with the strongest resource overlap (red box in Fig. 2(a)). As shown by Fig. 2(d), the time evolution of species abundances depends heavily on the environmental drive. Under fast environmental drive, the two species abundances track each other closely, while under slow drive, they exhibit a phase shift. When the environmental drive is mixed with fast and slow modes, the dynamics of each species also exhibit a mixture of modes, and the relation between the two species is less clear.
@@
Figure 2: The simulated guild based ecosystem. An example environment-mediated consumer-resource model (number of species $N=6$, number of resources $M=30$, and guild-structure bias $q=0.4$) demonstrating that the relation between the equal-time species abundance correlation $\mathbf{C}(0)$ and the underlying resource utilization overlap matrix $\mathbf{C}\mathbf{C}^{\intercal}$ depend on the timescale of the environmental drive. (a) Resource utilization overlap matrix $\mathbf{C}\mathbf{C}^{\intercal}$. The red boxes indicate an example pair of interacting species. (b) Environmental drives are chosen to be sinusoidal, with fast (red, $\omega_{\alpha}\in\left[1,10\right]$), slow (blue, $\omega_{\alpha}\in\left[0.01,0.1\right]$), and mixed timescales (green, $\omega_{\alpha}\in\left[0.03,3\right]$). (c) The correlation between the elements of equal-time correlation function $\mathbf{C}(0)$ and the elements of the resource-utilization matrix, $\mathbf{C}\mathbf{C}^{\intercal}$, is positive when the environmental drive is fast (red); negative when the drive is slow (blue), and shows no correlation when the drive is composed of a mixture of slow and fast timescales (green). In each panel, $r$ is Spearman’s rank correlation coefficient, $p$ is the p-value for the null hypothesis that there is no correlation between $\mathbf{C}(0)$ and $\mathbf{C}\mathbf{C}^{\intercal}$. Black arrows indicate the example pair of species in the red boxes of (a). (d) Example segments of time series of the log-transformed species abundances for the red-boxed pair of (a), under the three types of environmental drive given in panel (b). (e) Cross-correlation function between the species abundances of the example red-boxed pair of (a). Equal-time correlation is the value of $\mathbf{C}(\tau)$ at $\tau=0$.
@@ II.2.3 Spectral method for time-series analysis: Cross power spectral density (CPSD) and coherence
The maximum frequency is given by the Nyquist frequency, $f_{\text{max}}=1/2\Delta t_{\text{sampling}}$. The frequency resolution $\Delta f$ is chosen heuristically, such that the power spectral density estimate is relatively stable when $\Delta f$ is doubled or halved ($\Delta f=f_{\text{max}}/15$ for all the results presented in this manuscript). Using the simulated ECRM data, we estimate the power spectral density of the species abundance time series using the MVGC Toolbox mvgc_toolbox and the MATLAB built-in function with Welch’s method to reduce noise from using finite data. As shown in Fig. 3(b), compared to the equal-time correlation $\mathbf{C}(0)$ and its absolute value, both the total magnitude of CPSD $\mathcal{P}$ and the total coherence $\mathcal{TC}$ are much more informative of the underlying resource utilization overlap matrix $\mathbf{C}\mathbf{C}^{\intercal}$. Total coherence is the best of the four predictors. This advantage of the total coherence persists for both ECRMs under sinusoidal and OU drives and over the range of guild-structure biases $q$ (Fig. 3(c)).
@@ II.3 Binary classifier for guild-structure detection
Motivated by this, we asked whether total coherence and CPSD magnitude could be used to learn the large-scale guild structure of $\mathbf{C}\mathbf{C}^{\intercal}$. We used the four pairwise measures, $\mathbf{C}(0)$, $\lvert\mathbf{C}(0)\rvert$, $\mathcal{P}$, $\mathcal{TC}$, as scores to perform a binary classification between pairs of species belonging to the same guild and pairs belonging to different guilds. We applied this approach to our small $N=6$ simulated ECRM system with 2 guilds (Fig. 2(a)). To obtain a consistent partition into guilds, we perform single linkage clustering initiated from all pairs of species whose pairwise measure is larger than a threshold within the range of the pairwise measure. For example, if after thresholding species $i$ and $j$ are determined to belong to the same guild, as well as species $j$ and $k$, then all three species $i$, $j$, $k$ are grouped into to the same guild.
@@ II.4 Spectral methods work for general ECRM systems with larger sizes
In natural ecological systems, both the number of guilds and the number of species in each guild can be large. How well can the dynamical pairwise observables predict the resource utlization overlap for different system sizes? We now consider general ECRMs with larger number of guilds. To explore finite size scaling, we consider three families of ECRMs with numbers of species, resources, and guilds given in Table 1.
As we increase the size of the ECRMs by increasing the species size $N$, we also increase the number of resources in proportion, $M\propto N$, and the number of guilds as $k\propto\sqrt{N}$. We take $p_{\text{in}}=\left(1/2+q\right)\left(N/6\right)^{-1/2}$ and $p_{\text{out}}=\left(1/2-q\right)\left(N/6\right)^{-1}$. The scalings with $N$ are chosen such that the average connectivity degree for each species in the effective resource-utilization matrix $\mathbf{C}\mathbf{C}^{\intercal}$ becomes constant at large $N$ (see Methods IV.2 for details). The particular prefactors are chosen so that the numbers are consistent with the $N=6$ case considered earlier.
@@
We monitor the predictive performance of the same four pairwise measures as before ($\mathbf{C}(0)$, $\lvert\mathbf{C}(0)\rvert$, $\mathcal{P}$, and $\mathcal{TC}$) as we increase the system size from $N=6,k=2$ to $N=24,k=4$, and then to $N=96,k=8$ (Fig. 5(a,e,i)). We compute Spearman’s correlation coefficient between the pairwise observables and the elements of the $\mathbf{C}\mathbf{C}^{\intercal}$ matrix (Fig. 5(b, f, j)), and the ROC of guild detection after single linkage clustering (Fig. 5(d,h,i)).
@@
By the choice of scaling for $p_{\rm in}$ and $p_{\rm out}$, the $\mathbf{C}\mathbf{C}^{\intercal}$ matrix is sparse for all system sizes. We can set a threshold on our measures to perform link detection, i.e. identify whether there exists an effective interaction between species due to competition for the same resource. As shown in Fig. 5 for $q=0.5$ (and SI Fig. S1 for $q=0.4$), for all the three system sizes tested, the total coherence remains the best among the four predictors. The summary statistics in Fig. S2 show that the total coherence always performs the best among the four predictors, although all get worse as the system size increases. Guild detection (Fig. S2(c)) has a better AUC than link detection (Fig. S2(b)), especially for large system sizes, thanks to error correction afforded by single-linkage clustering.
@@
Figure 5: Pairwise observables for the simulated ECRM ecosystem of different sizes (Table 1), with $q=0.5$. (a) Effective resource-utilization matrix $\mathbf{C}\mathbf{C}^{\intercal}$ with $N=6$ species, $M=30$ resources, $k=2$ guilds, and $q=0.5$. (b) Correlation between pairwise measures of the species abundances and the resource utilization overlap. Measured pairwise observables include equal-time correlation $\mathbf{C}(0)$ and its absolute value $\lvert\mathbf{C}(0)\rvert$, the total absolute value of CPSD, and the total coherence. (c) ROC curve of link detection for the four pairwise observables. Same color code as panel (b); note that the total absolute value of CPSD overlaps with the total coherence, as both measures return perfect predictions. (d) Same as (c) but for guild detection, after performing single linkage clustering on the pairwise metrics with a tunable threshold. (e-h) Same as (a-d) but for $N=24$, $M=120$, and $k=4$. (i-l) Same as (a-d) but for $N=96$, $M=480$, and $k=8$. For all ECRMs, the environment is driven by OU processes with the intrinsic timescale set to $\omega_{\alpha}=1$. For all system sizes, the total duration of the simulated trajectory is set to $t_{f}=20000$, and the sampling timestep is set to $\Delta t_{\text{sampling}}=1$.
@@
II.5 Reconstruction of the $\mathbf{C}\mathbf{C}^{\intercal}$ matrix depends on sampling rate and total time series length
@@
The success of CPSD and coherence in detecting guild structures depends on the interplay of multiple timescales. These timescales include the doubling timescale $(\gamma r)^{-1}$ of the species, their typical lifetime $d_{x}^{-1}$, the environmental fluctuation timescales $\omega_{\alpha}^{-1}$, and their relaxation time $d_{R}^{-1}$. Additionally, there are experimental timescales: the total time $t_{f}$ of data acquisition and sampling interval $\Delta t_{\rm sampling}$. Here, we discuss the effects of the finite time series and sampling rate on recovering the elements and structure of the $\mathbf{C}\mathbf{C}^{\intercal}$ matrix.
@@
As we increase the total time of acquisition $t_{f}$, the accuracy of guild detection improves for all 4 observables as measured by AUC, as expected (Fig. 6(a)-(d)). For equal-time measurements ($\mathbf{C}(0)$ and $\lvert\mathbf{C}(0)\rvert$, (Fig. 6(a), (b)), the sampling interval $\Delta t_{\rm sampling}$ does not change the prediction significantly. However, for the total absolute CPSD and the total coherence (Fig. 6(c), (d)), the AUC for guild detection shows an optimal near $\Delta t_{\text{sampling}}=2$ and $\Delta t_{\text{sampling}}=0.5$ respectively, for which the AUC for guild detection is largest for the biggest range of total data acquisition timescale. For comparison, in our simulation the species death rate is $d_{x}=0.3$, corresponding to a turnover timescale of 3. In the limit of small $t_{f}$ and large $\Delta t_{\text{sampling}}$, the number of time points is less than the number of the time points required by spectral method given the frequency resolution we specified (for FFT the number of data points needs to be at least double the number of the frequency resolution), which means the methods cannot work (bottom right corner of Fig. 6(d)). Similar results hold for Spearman’s correlation between the pairwise metrics and the $\mathbf{C}\mathbf{C}^{\intercal}$ matrix element (SI Fig. S3).
@@
Figure 6: Influence of finite data duration and sampling rate on guild detection using pairwise measures of species abundance. AUC of the guild detection task using (a) Equal-time correlation $\mathbf{C}(0)$, (b) its absolute value $\lvert\mathbf{C}(0)\rvert$, (c) the total absolute value of CPSD, and (d) the total coherence, as the function of sampling duration $t_{f}$ and interval $\Delta t_{\rm sampling}$. For reference the species turnover timescale is $d_{x}^{-1}=3.33$. The ECRM is the same as the one considered in Fig. 3 with $N=6$, with OU environmental drive at $\omega_{\alpha}=1$ and guild-structure bias $q=0.4$.

### v2

@@ II.1 Environment-mediated consumer-resource model
Figure 1: Environment-mediated ecological interactions between resources and consumers. (a) Cartoon of interactions among environment-mediated resources and consumer species. (b) An example of the gain matrix for consumer-resource interaction $\mathbf{G}$ with elements $G_{i\alpha}=r_{i\alpha}\gamma_{i\alpha}$ (colorbar). The probability of non-zero consumer-resource interactions depends on whether the resource is preferred by the species or not, with an adjustable parameter $q$ that encodes guild-structure bias ($q=0.4$ for this example). All resources also have private resources to prevent extinction. (c) Effective interactions among the species are captured by the effective resource-utilization overlap matrix, $\mathbf{G}\mathbf{G}^{\intercal}$ (colorbar), shown here for 3 different values of the guild-structure bias $q$.
@@
The gain matrix for the consumer-resource interaction $\mathbf{G}$, where each entry is $G_{i\alpha}=r_{i\alpha}\gamma_{i\alpha}$, describes the intake of resource $\alpha$ by species $i$ (Fig. 1(b)). Summing over the resources we obtain the resource-utilization overlap matrix $\mathbf{G}\mathbf{G}^{\intercal}$ (Fig. 1(c)). Large off-diagonal entries of this matrix ($\mathbf{G}\mathbf{G}^{\intercal}_{i,j}$) indicate that species $i$ and $j$ have high overlap in the resources they utilize. Species with high resource-overlap are expected to compete, and those without are expected not to compete. One can consider the resource-utilization matrix as a proxy for effective interactions between species, which is informative about the community structure.
Specifically, we are interested in detecting guilds among the interacting species. To introduce a guild structure in resource overlap, we write the uptake rate as $r_{i\alpha}=rg_{i\alpha}$, where $g_{i\alpha}$ is an adjacency matrix that takes the value of 0 if species $i$ does not uptake resource $\alpha$, and 1 if it does consume this resource. In order to generate ensembles of random interaction strengths such that the resulting $\mathbf{G}\mathbf{G}^{\intercal}$ matrix has a block (guild) structure, we construct a bipartite graph between $N$ species and $M$ resources. We then separate both the species and the resources into $k$ groups, with each group of species preferring to consume resources from one group of resources. The elements of the adjacency matrix that specifies these preferences, $g_{i\alpha}$, are drawn from a Bernoulli distribution parametrized by a probability $p_{\text{in}}$. For the other $k-1$ groups of non-preferred resources, the elements of the adjacency matrix $g_{i\alpha}$ are drawn from a Bernoulli distribution parametrized by a probability $p_{\text{out}}<p_{\rm in}$. In addition, each species is assigned its private resource to prevent extinction (Fig. 1(b)). The yield $\gamma_{i\alpha}$ is drawn from a Gaussian distribution with a positive mean and small variance to introduce randomness. All yields are non-negative (see Methods for specific parameter values).
@@
The consumer-resource model (1)-(2), environmental drive (5)-(7), and resource guild structure $r_{i\alpha}$ among consumers, completely define the model. To illustrate its behavior and gain intuition, we first simulated a small network with $N=6$ species and $M=30$ resources, with a guild-structure bias of $q=0.4$. The resource-utilization matrix $\mathbf{G}\mathbf{G}^{\intercal}$ is given by Fig. 2(a). The specific set of parameters is given in Methods IV.1. The first $N_{a}=3$ species form one guild, while the remaining $N_{b}=3$ species form the other. To study the impact of environmental drive on species abundances, we simulate the ECRM under sinusoidal and OU environmental drives each sampled at three different ranges of timescales – slow, fast, and with a mixture of timescales (Fig. 2(b)). Example traces of pairs of species within the same guild show complex coupled dynamics depending on the environmental timescales (Fig. 2(d)).
@@ II.2 Pairwise measures of species couplings
To explore how the coupled dynamics of species abundances can be informative about their resource overlap $\mathbf{G}\mathbf{G}^{\intercal}$, we consider observables of increasing complexity that measure the deviation from independence between the time traces of two species.
@@ II.2.1 Equal-time correlation $\mathbf{C}(0)$
Specifically, as shown in Figure 2(c) in our example of 6 species distributed into 2 guilds, when the environmental drive is fast, $C_{ij}(0)\propto\left(\mathbf{G}\mathbf{G}^{\intercal}\right)_{ij}$, while when the environmental drive is slow, $C_{ij}(0)\propto-\left(\mathbf{G}\mathbf{G}^{\intercal}\right)_{ij}$. Intuitively this happens because fast resource dynamics drive coherent responses of members within a guild, but slow resource dynamics lead to competition within guilds (the underlying mechanism is first explored and explained in detail in crocker_microbial_2025 ). This problem is more pronounced when environmental fluctuations are driven by a mixture of timescales (middle panel in Fig. 2(c)). There is no correlation between the equal-time correlation $\mathbf{C}(0)$ and the resource utilization overlap $\mathbf{G}\mathbf{G}^{\intercal}$ ($p$-value = 0.4, Spearman correlation test). In this example, the pair of species with the strongest resource utilization overlap (indicated by red boxes in Fig. 2(a)) has a equal-time correlation close to 0, as indicated by the black arrow in Fig. 2(c).
Without knowing a priori the timescale of the environmental fluctuation, the equal-time correlation $\mathbf{C}(0)$ does not return a good enough estimator of the resource utilization overlap.
@@ II.2.2 Cross-correlation with time-delay reveals species interaction structure
Since equal-time correlations do not reveal the structure of effective species-species interactions, we took a closer look at time series data. Let us focus again on the pair of species with the strongest resource overlap (red box in Fig. 2(a)). As shown by Fig. 2(d), the time evolution of species abundances depends heavily on the environmental drive. Under fast environmental drive, the two species abundances track each other closely, while under slow drive, they exhibit a phase shift. When the environmental drive is mixed with fast and slow modes, the dynamics of each species also exhibit a mixture of modes, and the relation between the two species is less clear.
@@
Figure 2: The simulated guild based ecosystem. An example environment-mediated consumer-resource model (number of species $N=6$, number of resources $M=30$, and guild-structure bias $q=0.4$) demonstrating that the relation between the equal-time species abundance correlation $\mathbf{C}(0)$ and the underlying resource utilization overlap matrix $\mathbf{G}\mathbf{G}^{\intercal}$ depend on the timescale of the environmental drive. (a) Resource utilization overlap matrix $\mathbf{G}\mathbf{G}^{\intercal}$. The red boxes indicate an example pair of interacting species. (b) Environmental drives are chosen to be sinusoidal, with fast (red, $\omega_{\alpha}\in\left[1,10\right]$), slow (blue, $\omega_{\alpha}\in\left[0.01,0.1\right]$), and mixed timescales (green, $\omega_{\alpha}\in\left[0.03,3\right]$). (c) The correlation between the elements of equal-time correlation function $\mathbf{C}(0)$ and the elements of the resource-utilization matrix, $\mathbf{G}\mathbf{G}^{\intercal}$, is positive when the environmental drive is fast (red); negative when the drive is slow (blue), and shows no correlation when the drive is composed of a mixture of slow and fast timescales (green). In each panel, $r$ is Spearman’s rank correlation coefficient, $p$ is the p-value for the null hypothesis that there is no correlation between $\mathbf{C}(0)$ and $\mathbf{G}\mathbf{G}^{\intercal}$. Black arrows indicate the example pair of species in the red boxes of (a). (d) Example segments of time series of the log-transformed species abundances for the red-boxed pair of (a), under the three types of environmental drive given in panel (b). (e) Cross-correlation function between the species abundances of the example red-boxed pair of (a). Equal-time correlation is the value of $\mathbf{C}(\tau)$ at $\tau=0$.
@@ II.2.3 Spectral method for time-series analysis: Cross power spectral density (CPSD) and coherence
The maximum frequency is given by the Nyquist frequency, $f_{\text{max}}=1/2\Delta t_{\text{sampling}}$. The frequency resolution $\Delta f$ is chosen heuristically, such that the power spectral density estimate is relatively stable when $\Delta f$ is doubled or halved ($\Delta f=f_{\text{max}}/15$ for all the results presented in this manuscript). Using the simulated ECRM data, we estimate the power spectral density of the species abundance time series using the MVGC Toolbox mvgc_toolbox and the MATLAB built-in function with Welch’s method to reduce noise from using finite data. As shown in Fig. 3(b), compared to the equal-time correlation $\mathbf{C}(0)$ and its absolute value, both the total magnitude of CPSD $\mathcal{P}$ and the total coherence $\mathcal{TC}$ are much more informative of the underlying resource utilization overlap matrix $\mathbf{G}\mathbf{G}^{\intercal}$. Total coherence is the best of the four predictors. This advantage of the total coherence persists for both ECRMs under sinusoidal and OU drives and over the range of guild-structure biases $q$ (Fig. 3(c)).
@@ II.3 Binary classifier for guild-structure detection
Motivated by this, we asked whether total coherence and CPSD magnitude could be used to learn the large-scale guild structure of $\mathbf{G}\mathbf{G}^{\intercal}$. We used the four pairwise measures, $\mathbf{C}(0)$, $\lvert\mathbf{C}(0)\rvert$, $\mathcal{P}$, $\mathcal{TC}$, as scores to perform a binary classification between pairs of species belonging to the same guild and pairs belonging to different guilds. We applied this approach to our small $N=6$ simulated ECRM system with 2 guilds (Fig. 2(a)). To obtain a consistent partition into guilds, we perform single linkage clustering initiated from all pairs of species whose pairwise measure is larger than a threshold within the range of the pairwise measure. For example, if after thresholding species $i$ and $j$ are determined to belong to the same guild, as well as species $j$ and $k$, then all three species $i$, $j$, $k$ are grouped into to the same guild.
@@ II.4 Spectral methods work for general ECRM systems with larger sizes
In natural ecological systems, both the number of guilds and the number of species in each guild can be large. How well can the dynamical pairwise observables predict the resource utilization overlap for different system sizes? We now consider general ECRMs with larger number of guilds. To explore finite size scaling, we consider three families of ECRMs with numbers of species, resources, and guilds given in Table 1.
As we increase the size of the ECRMs by increasing the species size $N$, we also increase the number of resources in proportion, $M\propto N$, and the number of guilds as $k\propto\sqrt{N}$. We take $p_{\text{in}}=\left(1/2+q\right)\left(N/6\right)^{-1/2}$ and $p_{\text{out}}=\left(1/2-q\right)\left(N/6\right)^{-1}$. The scalings with $N$ are chosen such that the average connectivity degree for each species in the effective resource-utilization matrix $\mathbf{G}\mathbf{G}^{\intercal}$ becomes constant at large $N$ (see Methods IV.2 for details). The particular prefactors are chosen so that the numbers are consistent with the $N=6$ case considered earlier.
@@
We monitor the predictive performance of the same four pairwise measures as before ($\mathbf{C}(0)$, $\lvert\mathbf{C}(0)\rvert$, $\mathcal{P}$, and $\mathcal{TC}$) as we increase the system size from $N=6,k=2$ to $N=24,k=4$, and then to $N=96,k=8$ (Fig. 5(a,e,i)). We compute Spearman’s correlation coefficient between the pairwise observables and the elements of the $\mathbf{G}\mathbf{G}^{\intercal}$ matrix (Fig. 5(b, f, j)), and the ROC of guild detection after single linkage clustering (Fig. 5(d,h,i)).
@@
By the choice of scaling for $p_{\rm in}$ and $p_{\rm out}$, the $\mathbf{G}\mathbf{G}^{\intercal}$ matrix is sparse for all system sizes. We can set a threshold on our measures to perform link detection, i.e. identify whether there exists an effective interaction between species due to competition for the same resource. As shown in Fig. 5 for $q=0.5$ (and SI Fig. S1 for $q=0.4$), for all the three system sizes tested, the total coherence remains the best among the four predictors. The summary statistics in Fig. S2 show that the total coherence always performs the best among the four predictors, although all get worse as the system size increases. Guild detection (Fig. S2(c)) has a better AUC than link detection (Fig. S2(b)), especially for large system sizes, thanks to error correction afforded by single-linkage clustering.
@@
Figure 5: Pairwise observables for the simulated ECRM ecosystem of different sizes (Table 1), with $q=0.5$. (a) Effective resource-utilization matrix $\mathbf{G}\mathbf{G}^{\intercal}$ with $N=6$ species, $M=30$ resources, $k=2$ guilds, and $q=0.5$. (b) Correlation between pairwise measures of the species abundances and the resource utilization overlap. Measured pairwise observables include equal-time correlation $\mathbf{C}(0)$ and its absolute value $\lvert\mathbf{C}(0)\rvert$, the total absolute value of CPSD, and the total coherence. (c) ROC curve of link detection for the four pairwise observables. Same color code as panel (b); note that the total absolute value of CPSD overlaps with the total coherence, as both measures return perfect predictions. (d) Same as (c) but for guild detection, after performing single linkage clustering on the pairwise metrics with a tunable threshold. (e-h) Same as (a-d) but for $N=24$, $M=120$, and $k=4$. (i-l) Same as (a-d) but for $N=96$, $M=480$, and $k=8$. For all ECRMs, the environment is driven by OU processes with the intrinsic timescale set to $\omega_{\alpha}=1$. For all system sizes, the total duration of the simulated trajectory is set to $t_{f}=20000$, and the sampling timestep is set to $\Delta t_{\text{sampling}}=1$.
@@
II.5 Reconstruction of the $\mathbf{G}\mathbf{G}^{\intercal}$ matrix depends on sampling rate and total time series length
@@
The success of CPSD and coherence in detecting guild structures depends on the interplay of multiple timescales. These timescales include the doubling timescale $(\gamma r)^{-1}$ of the species, their typical lifetime $d_{x}^{-1}$, the environmental fluctuation timescales $\omega_{\alpha}^{-1}$, and their relaxation time $d_{R}^{-1}$. Additionally, there are experimental timescales: the total time $t_{f}$ of data acquisition and sampling interval $\Delta t_{\rm sampling}$. Here, we discuss the effects of the finite time series and sampling rate on recovering the elements and structure of the $\mathbf{G}\mathbf{G}^{\intercal}$ matrix.
@@
As we increase the total time of acquisition $t_{f}$, the accuracy of guild detection improves for all 4 observables as measured by AUC, as expected (Fig. 6(a)-(d)). For equal-time measurements ($\mathbf{C}(0)$ and $\lvert\mathbf{C}(0)\rvert$, (Fig. 6(a), (b)), the sampling interval $\Delta t_{\rm sampling}$ does not change the prediction significantly. However, for the total absolute CPSD and the total coherence (Fig. 6(c), (d)), the AUC for guild detection shows an optimal near $\Delta t_{\text{sampling}}=2$ and $\Delta t_{\text{sampling}}=0.5$ respectively, for which the AUC for guild detection is largest for the biggest range of total data acquisition timescale. For comparison, in our simulation the species death rate is $d_{x}=0.3$, corresponding to a turnover timescale of 3. In the limit of small $t_{f}$ and large $\Delta t_{\text{sampling}}$, the number of time points is less than the number of the time points required by spectral method given the frequency resolution we specified (for FFT the number of data points needs to be at least double the number of the frequency resolution), which means the methods cannot work (bottom right corner of Fig. 6(d)). Similar results hold for Spearman’s correlation between the pairwise metrics and the $\mathbf{G}\mathbf{G}^{\intercal}$ matrix element (SI Fig. S3).
@@
Figure 6: Influence of finite data duration and sampling rate on guild detection using pairwise measures of species abundance. AUC of the guild detection task using (a) Equal-time correlation $\mathbf{C}(0)$, (b) its absolute value $\lvert\mathbf{C}(0)\rvert$, (c) the total absolute value of CPSD, and (d) the total coherence, as the function of sampling duration $t_{f}$ and interval $\Delta t_{\rm sampling}$. For reference the species turnover timescale is $d_{x}^{-1}=3.33$. The results are averaged over 50 random realizations of the ECRMs with $N=6$, with OU environmental drive at $\omega_{\alpha}=1$ and guild-structure bias $q=0.4$.